OpenStudy (anonymous):

Solve for x: |2x + 6| − 4 = 20

OpenStudy (anonymous):

2x+6-4=20 2x+2=20 2x=18 x=9

hero (hero):

Hint: 1. Add 4 to both sides: |2x + 6| = 20 + 4 2. Simplify right side: |2x + 6| = 24 3. Inverse absolute value of both sides: 2x + 6 = ±24 4. Split into two equations: 2x + 6 = 24 2x + 6 = -24 5. Solve each equation separately: 2x = 24 -6 2x = -24 -6 2x = 18 2x = -30 x = 18/2 x = -30/2 x = 9 x = -15

hero (hero):

Any questions?